# A longitudinal Bayesian framework for estimating causal dose-response relationships

### v2

@@ 3.2 Example 2
Similar to the first example, we fit a Poisson GEE model for the outcome of $Y_{i}$ with the inclusion of covariates $D_{1it}$, the three-order spline of estimated GPS for COV method, and as for WOR method, we fit the outcome mean with the three-order spline of $D_{1it}$ with the GPS as the weight. For the DP approach, we use the same steps to sample the data with size 500, but the outcome $y^{s}_{i}$ then is sample from a Poisson distribution with mean $g(\mathbb{E}[Y_{i}\mid D_{i},\bar{X}_{i},\widehat{e}_{i}(\bar{X}_{i})])$, which is the fitted value from a GEE model with $y^{s^{\prime}}_{i}$ with the inclusion of covariates $d^{s}_{i}$, the three-order spline of estimated GPS.
Figure 3 demonstrates the 1000 posterior dose-response curves with observed data points. We observe a consistent pattern where the weighting method leads to greater variance, particularly in regions with fewer observations. Similar to the previous example, there is not much difference between the BB and DP approaches. Table 1 summarizes the results of 1,000 simulation runs at three specified dose levels. The COV and WOR methods both give roughly unbiased estimates for all dose levels. The WOR approach consistently exhibits higher average estimated variances compared to the COV method. When applying the weighting approach, the DP method continues to demonstrate slightly lower variability relative to the BB method. This is likely because the DP resampling scheme yields larger sample sizes, which help stabilize estimates, particularly in regions with sparse data. In terms of coverage performance, the BB-COV procedure performs well at dose level 4, where data are more abundant. However, at other dose levels, where observations are sparse, the coverage rate for the COV method drops substantially. This suggests that the resulting credible intervals may be too narrow in regions with limited data support, leading to undercoverage. On the other hand, the WOR-BB procedure maintains a relatively stable undercoverage rate, ranging from approximately 82 to 85 across all dose levels. However, the DP method consistently achieves coverage around the nominal target across all three dose levels, regardless of whether it is used with the COV or WOR estimating procedure. This result demonstrates that the DP approach better accommodates data sparsity and improves posterior coverage across the dose-response curve.
Figure 3: Example 2: Posterior dose-response curves using WOR and COV approaches for the Poisson outcome.
@@ 4 Application: Quantifying the causal impact of mass transit on COVID-19 transmission
As discussed in the introduction, empirical studies quantifying the causal impact of mass public transport on COVID-19 transmission are scarce. Most use aggregate measures of mobility from mobile phone data sources such as those that were made publicly available by Google Mobility and Apple Maps during the pandemic, and issues of causality have received little attention. In our review, we have identified just one empirical study by Romero García et al. (2023) which uses causal analysis methods to assess the impact of travel specifically on metro networks (referred to as subways in the paper) on the transmission of COVID-19 in Valencia, Spain. In the following sections, we provide a summary of empirical studies undertaken on the impact of mobility on COVID-19 transmission, beginning with an overview of literature using non-causal analysis methods, followed by a summary of the limited literature adopting causal methods.
@@ 4.1 Non-causal empirical studies
There is a vast number of empirical studies that use non-causal statistical analysis methods to study the association of mobility with COVID-19 transmission as measured by either the number of COVID-19 cases or the virus reproduction number. Generally, most studies conclude that an increase in mobility is positively associated with an increase in COVID-19 transmission, and restrictions on mobility are associated with reductions in COVID-19 transmission. Studies using statistical correlation methods are prevalent in the literature. For example, Jia et al. (2020) reported a positive linear correlation between mobility and case incidence in China at the early stages of the pandemic. Badr et al. (2020) analysed mobility patterns in the USA and also reported positive linear correlations between mobility and COVID-19 case rates. Kissler et al. (2020) used Pearson’s linear correlation coefficient and reported that their estimates of COVID-19 prevalence were negatively associated with reductions in mobility in their analysis of New York City.
Studies adopting different forms of multivariate regression methods are also prevalent, and here we present a selection of examples. Li et al. (2021) used multivariate linear regression and reported that the reproduction number of the virus increased with higher levels of mobility at retail and recreation places, workplaces, and transit stations in the UK. Kraemer et al. (2020) applied a generalized linear regression model framework and found that mobility was positively correlated with case incidence in China prior to travel restrictions, but the relationship was no longer statistically significant after the implementation of restrictions. McGrail et al. (2020) undertook a cross-country analysis of 134 countries and 50 US states. The authors used a generalized linear mixed effects model and reported that mobility and infection rates reduced after the imposition of social distancing restrictions. Islam et al. (2020) also performed a cross-country analysis of 149 countries using a two-stage linear regression analysis. It was reported that case incidence reduced with the imposition of public transport closures, school closures, work closures, restrictions on mass gatherings, and blanket lockdowns. Xiong et al. (2020) reported a positive relationship between mobility and COVID-19 cases in the US using structural equation modeling with dynamic panel and time-varying effects at the level of US counties. Manzira et al. (2022) used a multivariate time series regression model with autoregressive lag effects and reported a positive association between mobility and the incidence of COVID-19 in the city centre areas of Dublin, Ireland.
@@ 4.2 Causal empirical studies
Steiger et al. (2021), Nugent and Balzer (2023), Cho et al. (2022), and Chernozhukov et al. (2021) used aggregate indicators of mobility from mobile phone data records unattributed to specific travel modes to assess causality between mobility and COVID-19 cases. Steiger et al. (2021) estimated the causal impact of factors affecting the number of new cases of COVID-19 in Germany using a DAG and a negative binomial regression model. The authors found a positive causal impact between mobility in retail and recreational spaces and workplaces and COVID-19 cases, and a negative causal impact for essential shopping and mobility within residential areas. Other influential factors included weather, socio-demographic characteristics, and public awareness of COVID-19. Nugent and Balzer (2023) adopted the causal modified treatment policy method with the Super Learner machine learning algorithm to quantify the impact of mobility on COVID-19 cases in US counties. The authors found that unadjusted estimates suggested that decreasing mobility is causally linked to decreasing COVID-19 cases, but found that their adjusted estimates gave mixed results indicating potential confounding in the unadjusted estimates, which were most likely attributed to unknown differences in county characteristics. Cho et al. (2022) also analysed the causal relationship between COVID-19 cases and mobility in the US. They adopted an ensemble empirical mode decomposition method with causal decomposition and reported that mobility levels have a causal relationship with long-term variations in COVID-19 cases rather short-term variations. Chernozhukov et al. (2021) used a causal structural model framework to test a range of counterfactual experiments including mobility changes and their impact on COVID-19 cases in the US. The authors find that reductions in mobility due to stay at home orders and business closures were effective in reducing case numbers, while the impacts of reductions in mobility due to school closures were inconclusive.
@@ 4.3 Data
The Community of Metros (COMET) benchmarking group administered by the Transport Strategy Centre at Imperial College London have provided ridership data from 8 international member metro networks. In keeping with commercial confidentiality requirements, the metro identities have been anonymized and are referred to with a numerical index ranging from 1 to 8 throughout this paper. The ridership data reports the numerical ridership figures for each day. The data are recorded from before the onset of the COVID-19 pandemic to early 2022, with the earliest and latest dates in the data set being 1 March 2020 and 23 January 2022.
@@
For context, a summary of all variables including their descriptive statistics aggregated across all metros in the analysis is given in Table 2. The ridership is transformed in a logarithmic scale, cases and deaths variables are transformed with $\log(x+1)$ scale. From our preliminary analysis, the distribution of ridership is highly skewed and therefore we take the right $\log$ scale. After the transformation, the distribution looks more normal, and the ridership in $\log$ scale will be the treatment variable. For the outcome variable, we will consider to use cases without transformation and in $\log(x+1)$ scale.
@@ 4.4 Results
To estimate the dose-response relationship between the ridership and COVID-19 cases, we employed the proposed Bayesian method to analyze two different outcomes: the raw case counts, modeled using a log link with the Poisson mean-variance relationship, and the log-transformed case counts, $log(x+1)$, modeled under an identity link with the Gaussian mean-variance relationship. For both outcomes, we specified an independent working correlation structure and included all relevant confounders along with a cubic spline function of the GPS as covariates. Similar to the simulation, for the GPS model, we used a GEE model that adjusts for all time-varying confounders. We generated a total of 5,000 posterior samples. For the DP method, we generate a synthetic dataset of size $J=200$ using a procedure similar to that employed in the simulation studies. Due to the relatively small sample size, the stick-breaking weights in the DP decreased rapidly, falling below $10^{-8}$ after approximately 200 iterations. Figure 4 shows the posterior predictive dose-response curves for WOR and COV methods. When using the COV method, we observe an increasing dose-response trend for both outcomes, suggesting that higher metro ridership is associated with a greater rise in COVID-19 cases. However, the variability in the estimates is higher in certain dose ranges (e.g, 10-12 and 14–15). The DP approach shows a more stable curve as the synthetic dataset have a larger sample. When using case counts as the outcome, the COV method shows a generally exponential increase in cases as the logarithm of ridership rises. In contrast, the DP method appears to overcompensate in certain regions, resulting in some non-smooth curves likely because some portion of the resampled data is drawn from areas that deviate from the main trend. The WOR method, on the other hand, shows an initial rise in case counts that eventually plateaus and then slightly declines. In regions with sparse data, the DP approach produces more stable estimates, suggesting improved performance in low-density areas. We also report the estimated APOs at various doses of the posterior sample in Table 3. We choose the dose level at 11, 13 and 15, which corresponds to 2.5%, 50% and 97.5% quantiles of the logarithm of ridership, respectively. Similar to the observation in Figure 4, the COV method reveals that the number of cases increases as the logarithm of ridership rises. On average, the number of cases increases from about 7330 to 59000 when dose level changes from 11 to 15, using $\log$ (case +1+1) as the outcome. When modeling the outcome as the count data, the number of cases increases by an even higher extent, from approximately 500000 to 5000000. This increase is attributed to the fact that exponential change when modeling the count data. However, the WOR method shows a different trend. While the number of cases initially increases, it begins to decline as the dose level rises from 13 to 15. In summary, the number of cases tends to increase as the dose level—representing public transport usage—rises, but this relationship is not strictly linear. While both the log-transformed and count-based outcome suggest a strong upward trend in case numbers at lower dose regions, the log-transformed model shows a more moderate increase due to its smoothing effect, whereas the count model captures a steeper, exponential rise.
Figure 4: Application: Posterior dose-response curves using WOR (right panel) and COV (left panel) approaches WOR and COV approaches, with outcomes defined as $\log$(case +1+1) and cases.
Table 3: Posterior estimates of the dose-response function and associated APOs with credible intervals.
Outcome: $\log$(case +1+1)
Method dose = 11 dose = 13 dose = 15
COV-$B$ Posterior median 8.83 10.32 10.86
Credible interval (7.98, 9.76) (9.72, 10.76) (9.39, 12.40)
WOR-$B$ Posterior median 8.85 10.35 9.95
Credible interval (8.00, 9.95) (9.49, 10.71) (8.86, 11.63)
COV-DP Posterior median 8.90 10.30 11.11
Credible interval (8.30, 9.56) (9.91, 10.30) (10.06, 12.15)
WOR-DP Posterior median 8.98 10.39 10.31
Credible interval (8.28, 9.72) (9.93, 10.76) (8.92, 11.81)
Outcome: Cases
COV-$B$ Posterior median $5.90\times 10^{5}$ $1.63\times 10^{6}$ $5.41\times 10^{6}$
Credible interval $(2.07\times 10^{5},1.2\times 10^{6})$ $(1.29\times 10^{6},1.90\times 10^{6})$ $(2.22\times 10^{6},1.13\times 10^{7})$
WOR-$B$ Posterior median $1.19\times 10^{6}$ $1.90\times 10^{6}$ $1.52\times 10^{6}$
Credible interval $(2.51\times 10^{5},4.03\times 10^{6})$ $(1.46\times 10^{6},3.24\times 10^{6})$ $(4.49\times 10^{5},2.38\times 10^{7})$
COV-DP Posterior median $4.17\times 10^{5}$ $1.05\times 10^{6}$ $1.69\times 10^{6}$
Credible interval $(0.41\times 10^{5},2.26\times 10^{6})$ $(3.85\times 10^{5},3.34\times 10^{6})$ $(1.00\times 10^{5},1.85\times 10^{7})$
WOR-DP Posterior median $4.94\times 10^{5}$ $1.25\times 10^{6}$ $9.41\times 10^{5}$
Credible interval $(0.69\times 10^{5},1.66\times 10^{6})$ $(5.68\times 10^{5},2.71\times 10^{6})$ $(0.90\times 10^{5},7.10\times 10^{6})$
@@ 5 Discussion
In this paper, we introduced a non-parametric Bayesian framework for estimating longitudinal causal dose-response relationships in the presence of repeated measures. We address the issue of Bayesian inference in the context of model misspecification, and tackle this problem with a constraint optimization where the approximated model needs to satisfy some moment conditions while minimizing some distance between the true data generating model. From this route, we argue that the general Bayesian bootstrap approach with the DP specification is the consequence of the prior on the non-parametric probabilities. Our approach extends traditional Bayesian methodology by incorporating the double non-parametric generalized Bayesian bootstrap, while explicitly capturing the temporal structure of the data through the GEE framework. We model both GPS and the outcome mean via the GEE, and incorporate the GPS into the outcome model in two ways, as a covariate and through weighting. From the simulation studies, both of the resulting Bayesian dose-response estimators exhibit approximately unbiased estimation and achieve nominal posterior coverage rates under a range of scenarios using the DP method. However, the weighting-based estimator tends to have a larger posterior variance and displays greater instability when few observations are available at specific levels of the exposure. This finding corresponds to other similar inverse treatment weighting approaches with continuous exposure where binning of the continuous exposure may be considered (Naimi et al., 2014). In our empirical study, we investigated the impact of metro ridership on COVID-19 case counts across major urban centers during the pandemic. The results suggest that increased metro ridership contributes to the spread of COVID-19 up to a certain exposure threshold, beyond which the effect plateaus. These findings have important implications for pandemic-related urban policy, particularly in informing lockdown strategies and optimizing transportation systems in future public health crises.
@@
One appealing aspect of Bayesian estimators is their flexibility, as they allow for flexible approaches with minimal assumptions about the dose-response relationship, and for the prior knowledge to be added in the model. The proposed method can easily handle conditionally censored data by incorporating time-varying censoring weights, and moreover, it allows for the inclusion of prior information via the sampling-importance resampling approach (Saarela et al., 2015). Bayesian methods are increasingly influential in applied causal inference, offering the ability to make direct probabilistic statements about treatment effects and to assess sensitivity with respect to alternative priors or expert-driven inputs. Despite these advantages, Bayesian causal methods for longitudinal data remain underdeveloped Liu et al. (2020). In this paper, we proposed a scalable Bayesian approach that effectively handles longitudinal data and is easily implementable. Our proposed Bayesian dose-response estimation is derived under the SUTVA assumption, assuming no interference. In our motivating COVID study, this assumption translates to each unit’s monthly metro-ridership exposure being assumed to affect only its own subsequent COVID-19 risk, which can be violated in application. Methodological extensions that relax SUTVA are emerging (O’Riordan and Gilligan-Lee, 2025). Future work will adapt these ideas to longitudinal, continuous-dose settings, enabling joint estimation of direct and indirect (spillover) ridership effects on COVID-19 dynamics. Despite this limitation, the proposed estimator fills an important methodological gap by coupling flexible non-parametric Bayesian inference with repeated-measures data and a time-varying continuous exposure. It remains fully appropriate for applications where interference is unlikely (e.g., individual risk-factor studies in chronic-disease epidemiology). Furthermore, our framework offers a robust foundation that broader spillover-aware extensions can build upon.
Acknowledgements

### v3

@@ 3.2 Example 2
Similar to the first example, for the GEE associated with the first stage of GPS estimation, we include only the covariates $X_{1it}$ and $X_{2it}$. For the outcome model, we fit a Poisson GEE for $Y_{i}$. In the COV method, the model includes $D_{1it}$ and a cubic spline of the estimated GPS. In the WOR method, the outcome mean is modeled using a cubic spline of $D_{1it}$, with the estimated GPS incorporated as a weight. For the DP-based method, the predicted mean is obtained from a GEE outcome model including $D_{it}$ and a cubic spline of the estimated GPS.
Figure 3 demonstrates the 1000 posterior dose-response curves with observed data points. Both BB-based and DP-based approaches closely align with the true dose–response curve, with only small deviations at the high-dose tail. The WOR-based formulations exhibit wider 95% credible intervals, reflecting increased uncertainty in regions with limited data support. Similar with the previous example, the DP-based methods produce slightly wider intervals at extreme doses, capturing additional uncertainty induced by the more flexible prior. The Bayes-MSM method shows the most deviation of the posterior mean from the true curve at high doses, suggesting that MSM can lead to certain bias in regions where data are sparse.
Table 1 summarizes the results of 1,000 simulation runs at three specified dose levels. All methods both give roughly unbiased estimates for all dose levels. WOR-based methods again exhibit larger standard deviations, with WOR-DP achieving near-nominal coverage, while WOR-BB shows systematic undercoverage across all doses. The COV-based methods maintain good coverage at all dose levels. Similar to the previous example, Bayes-MSM shows undercoverage at the dose extremes, highlighting its sensitivity in regions with limited data support as we observed in Figure 3.
Figure 3: Example 2: Posterior predictive dose–response curves for the Poisson outcome estimated using the Bayesian bootstrap, Dirichlet process, and Bayesian parametric MSM approaches.
@@ 4 Application: Quantifying the causal impact of mass transit on COVID-19 transmission
As discussed in the introduction, empirical studies quantifying the causal impact of mass public transport on COVID-19 transmission are scarce. For estimation, we have monthly (longitudinal) city level data available which measure Covid cases, mass transit ridership, mobility containment measures, and a set of mobility characteristics for the period between 1 March 2020 and 23 January 2022. Longitudinal data provide the temporal resolution needed to understand how public transit ridership and COVID-19 incidence evolve and intersect. Most use aggregate measures of mobility from mobile phone data sources such as those that were made publicly available by Google Mobility and Apple Maps during the pandemic, and issues of causality have received little attention. In our review, we have identified just one empirical study by Romero García et al. (2023) which uses causal analysis methods to assess the impact of travel specifically on metro networks (referred to as subways in the paper) on the transmission of COVID-19 in Valencia, Spain. In the following sections, we provide a summary of empirical studies undertaken on the impact of mobility on COVID-19 transmission, beginning with an overview of literature using non-causal analysis methods, followed by a summary of the limited literature adopting causal methods.
@@ 4.1 Non-causal empirical studies
There is a vast number of empirical studies that use non-causal statistical analysis methods to study the association of mobility with COVID-19 transmission as measured by either the number of COVID-19 cases or the virus reproduction number. Generally, most studies conclude that an increase in mobility is positively associated with an increase in COVID-19 transmission, and restrictions on mobility are associated with reductions in COVID-19 transmission. Studies using statistical correlation methods are prevalent in the literature. For example, Jia et al. (2020) reported a positive linear correlation between mobility and case incidence in China at the early stages of the pandemic. Badr et al. (2020) analyzed mobility patterns in the USA and also reported positive linear correlations between mobility and COVID-19 case rates. Kissler et al. (2020) used Pearson’s linear correlation coefficient and reported that their estimates of COVID-19 prevalence were negatively associated with reductions in mobility in their analysis of New York City.
Studies adopting different forms of multivariate regression methods are also prevalent, and here we present a selection of examples. Li et al. (2021) used multivariate linear regression and reported that the reproduction number of the virus increased with higher levels of mobility at retail and recreation places, workplaces, and transit stations in the UK. Kraemer et al. (2020) applied a generalized linear regression model framework and found that mobility was positively correlated with case incidence in China prior to travel restrictions, but the relationship was no longer statistically significant after the implementation of restrictions. McGrail et al. (2020) undertook a cross-country analysis of 134 countries and 50 US states. The authors used a generalized linear mixed effects model and reported that mobility and infection rates reduced after the imposition of social distancing restrictions. Islam et al. (2020) also performed a cross-country analysis of 149 countries using a two-stage linear regression analysis. It was reported that case incidence reduced with the imposition of public transport closures, school closures, work closures, restrictions on mass gatherings, and blanket lockdowns. Xiong et al. (2020) reported a positive relationship between mobility and COVID-19 cases in the US using structural equation modeling with dynamic panel and time-varying effects at the level of US counties. Manzira et al. (2022) used a multivariate time series regression model with autoregressive lag effects and reported a positive association between mobility and the incidence of COVID-19 in the city center areas of Dublin, Ireland.
@@ 4.2 Causal empirical studies
Steiger et al. (2021), Nugent and Balzer (2023), Cho et al. (2022), and Chernozhukov et al. (2021) used aggregate indicators of mobility from mobile phone data records unattributed to specific travel modes to assess causality between mobility and COVID-19 cases. Steiger et al. (2021) estimated the causal impact of factors affecting the number of new cases of COVID-19 in Germany using a DAG and a negative binomial regression model. The authors found a positive causal impact between mobility in retail and recreational spaces and workplaces and COVID-19 cases, and a negative causal impact for essential shopping and mobility within residential areas. Other influential factors included weather, socio-demographic characteristics, and public awareness of COVID-19. Nugent and Balzer (2023) adopted the causal modified treatment policy method with the Super Learner machine learning algorithm to quantify the impact of mobility on COVID-19 cases in US counties. The authors found that unadjusted estimates suggested that decreasing mobility is causally linked to decreasing COVID-19 cases, but found that their adjusted estimates gave mixed results indicating potential confounding in the unadjusted estimates, which were most likely attributed to unknown differences in county characteristics. Cho et al. (2022) also analyzed the causal relationship between COVID-19 cases and mobility in the US. They adopted an ensemble empirical mode decomposition method with causal decomposition and reported that mobility levels have a causal relationship with long-term variations in COVID-19 cases rather short-term variations. Chernozhukov et al. (2021) used a causal structural model framework to test a range of counterfactual experiments including mobility changes and their impact on COVID-19 cases in the US. The authors find that reductions in mobility due to stay at home orders and business closures were effective in reducing case numbers, while the impacts of reductions in mobility due to school closures were inconclusive.
@@ 4.3 Data
The Community of Metros (COMET) benchmarking group administered by the Transport Strategy Centre at Imperial College London have provided ridership data from 8 international member metro networks. In keeping with commercial confidentiality requirements, the metro identities have been anonymized and are referred to with a numerical index ranging from 1 to 8 throughout this paper. hlTherefore, the study units correspond to eight distinct geographical locations. The ridership data reports the numerical ridership figures for each day. The data are recorded from before the onset of the COVID-19 pandemic to early 2022, with the earliest and latest dates in the data set being 1 March 2020 and 23 January 2022.
@@
For context, a summary of all variables including their descriptive statistics aggregated across all metros in the analysis is given in Table 2. The ridership is transformed in a logarithmic scale, and the number of cases are transformed with $\log(x+1)$ scale. From our preliminary analysis, the distribution of ridership is highly skewed and therefore we take the right $\log$ scale. After the transformation, the distribution looks more normal, and the ridership in $\log$ scale will be the treatment variable. For the outcome variable, we will consider to use cases without transformation and in $\log(x+1)$ scale.
@@ 4.4 Results
To estimate the dose-response relationship between the ridership and COVID-19 cases, we employed the proposed Bayesian method to analyze two different outcomes: the raw case counts, modeled using a log link with the Poisson GEE, and the log-transformed case counts, $log(x+1)$, modeled under an identity link with the Gaussian GEE. For both outcomes, we specified an independent working correlation structure and included all relevant confounders along with a cubic spline function of the GPS as covariates. Similar to the simulation, for the GPS model, we used a GEE model that adjusts for all time-varying confounders. We generated a total of 5,000 posterior samples. For the DP method, we generate a synthetic dataset of size $N=200$ using a procedure similar to that employed in the simulation studies. Due to the relatively small sample size, the stick-breaking weights in the DP decreased rapidly, falling below $10^{-8}$ after approximately 200 iterations. Figure 4 presents posterior predictive dose–response curves estimated using COV and WOR methods. Across all specifications, the estimated dose–response relationship indicates an overall positive and monotone causal effect of mobility on COVID-19 cases, with higher ridership associated with higher expected case counts. Under the COV-based approach, the estimated dose–response function increases for both log-transformed and count outcomes, although uncertainty is larger in certain dose ranges (notably between 10–12 and 14–15) due to reduced data density. The DP-based approach shows a more stable curve as the synthetic dataset have a larger sample. When using case counts as the outcome, the COV-based method shows a generally exponential increase in cases as the logarithm of ridership rises. However, the DP-based methods appear to overcompensate in certain regions, resulting in some non-smooth curves likely because some portion of the resampled data is drawn from areas that deviate from the main trend. In contrast, the WOR estimator shows a rapid increase in cases at lower dose levels followed by a plateau and slight decline at higher doses, indicating sensitivity to reweighting in sparsely populated regions. In regions with sparse data, the DP-based approach produces more stable estimates, suggesting improved performance in low-density areas. In summary, the number of cases tends to increase as the dose level, representing that public transport usage rises, but this relationship is not strictly linear. While both the log-transformed and count-based outcome suggest a strong upward trend in case numbers at lower dose regions, the log-transformed model shows a more moderate increase due to its smoothing effect, whereas the count model captures a steeper, non-smooth and exponential rise.
Figure 4: Application: Posterior predictive for dose–response estimates for ridership ($\log$ scale) and COVID-19 cases. Panels labeled $\log(x+1)$ use Gaussian GEE models for log-transformed cases, while panels labeled $\log$ scale use Poisson GEE models with posterior predictive samples rescaled to the $\log$ scale. Results are shown for Bayesian bootstrap and DP inference under covariate-adjusted and weighted outcome regression specifications.
@@ 5 Discussion
In this paper, we introduced a non-parametric Bayesian framework for estimating longitudinal causal dose-response relationships in the presence of repeated measures. The proposed framework addresses Bayesian inference under model misspecification by adopting a GEE perspective. Instead of specifying a fully parametric likelihood, the target estimand is defined through moment conditions that characterize the marginal causal parameters of interest, and uncertainty is quantified by placing a nonparametric prior on the underlying data-generating distribution. Specifically, this approach extends traditional Bayesian methodology by integrating the generalized Bayesian bootstrap within a GEE framework, thereby explicitly accounting for the temporal structure of longitudinal data. Both the GPS and the outcome mean are modeled using GEEs, and the GPS is incorporated into the outcome model in two ways: as an additional covariate or through stabilized inverse weighting. Simulation results indicate that both Bayesian dose–response estimators are approximately unbiased and achieve near-nominal posterior coverage across a range of scenarios when using the DP specification. However, the weighting-based estimator exhibits larger posterior variance, a behavior consistent with other inverse-weighting approaches for continuous treatments, where exposure binning is often required to mitigate variability (Naimi et al., 2014). In our empirical study, we investigated the impact of metro ridership on COVID-19 case counts across major urban centers during the pandemic. The results suggest that increased metro ridership contributes to the spread of COVID-19 up to a certain exposure threshold, beyond which the effect plateaus. These findings have important implications for pandemic-related urban policy, particularly in informing lockdown strategies and optimizing transportation systems in future public health crises.
@@
One appealing aspect of Bayesian estimators is their flexibility, as they allow for flexible approaches with minimal assumptions about the dose-response relationship, and for the prior knowledge to be added in the model. The proposed method can easily handle conditionally censored data by incorporating time-varying censoring weights. Bayesian methods are increasingly influential in applied causal inference, offering the ability to make direct probabilistic statements about treatment effects and to assess sensitivity with respect to alternative priors or expert-driven inputs. Despite these advantages, Bayesian causal methods for longitudinal data remain underdeveloped (Liu et al., 2020). In this paper, we proposed a scalable Bayesian approach that effectively handles longitudinal data and is easily implementable. Our proposed Bayesian dose-response estimation is derived under the SUTVA assumption, assuming no interference. In our motivating COVID study, this assumption translates to each unit’s monthly metro-ridership exposure being assumed to affect only its own subsequent COVID-19 risk, which can be violated in application. Methodological extensions that relax SUTVA are emerging (O’Riordan and Gilligan-Lee, 2025). Future work will adapt these ideas to longitudinal, continuous-dose settings, enabling joint estimation of direct and indirect (spillover) ridership effects on COVID-19 dynamics. Despite this limitation, the proposed estimator fills an important methodological gap by coupling non-parametric Bayesian inference with repeated-measures data and a time-varying continuous exposure. It remains fully appropriate for applications where interference is unlikely (e.g., individual risk-factor studies in chronic-disease epidemiology). Furthermore, our framework offers a robust foundation that broader spillover-aware extensions can build upon.
Acknowledgement